OpenStudy (anonymous):

-1/2x=-7/8 use multiplication principle x=?

OpenStudy (anonymous):

The final answer will be: x=7/4

OpenStudy (anonymous):

x=7/4

OpenStudy (anonymous):

multiplication principle means if a=b, then a*c=b*c, so if you multiply both sides by -2, you get your answer

OpenStudy (anonymous):

It's 7/4